OpenStudy (anonymous):

Find the value of y when x equals 1 in this equation. 2x + (-7y) = -54

OpenStudy (anonymous):

+ -7y = -7y, right?

OpenStudy (anonymous):

so just substitute x with 1; 2(1)+(-7y)=-54 Then just do the equation: 2-7y=-54 arrange the unknowns and numbers: -7y=-54-2 and just solve: -7y=-56 y=-56/-7 And the answer: y=8

OpenStudy (anonymous):

Thank you, the explanation will help with my other problems too!

OpenStudy (anonymous):

heehees no probs :) im glad i helped you ^^ good luck on your other questions ;D